In the study of fingerprints, an important quantitative characteristic is the total ridge count for 10 fingers of an individual. Suppose that the total ridge counts of individuals in a certain population are approximately normally distributed with a mean of 140 and a standard deviation of 50. Find the probability that an individual picked at random from this population will have a ridge count of between 100 and 200.
A) .8849
B) .6730
C) .7881
E) .7875

Answer by Boreal(15235)

for 100, z is (100-140)/50=-4/5
for 200, z is (200-140)/50=6/5
the probability of something between z of -0.8 and z of 1.2 is 0.6730 B.